OpenStudy (anonymous):

Find the limit as x approaches negative infinity of (6t^2 + 5t)/(1-t)(2t-3) Step by step would be great please!

OpenStudy (debbieg):

Do you mean as \[t \rightarrow-\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\lim_{t \rightarrow -\infty} \] typing problem

OpenStudy (debbieg):

\[\lim_{t \rightarrow -\infty}\frac{ 6t^2+5t }{ (1-t)(2t-3) }=\lim_{t \rightarrow -\infty}\frac{ 6t^2+5t }{ -2t^2+5t-3 }=\lim_{t \rightarrow -\infty}\frac{ 6t^2 }{ -2t^2 }=-3\] The ratio of the coefficients of the leading terms determine the limits at infinity of a rational function when the degree of the numerator = degree of denominator.